# ColorVideoVDP: A visual difference predictor for image, video and display distortions

### v1

@@ 3.1 Display model
$n_{\text{ppd}}=\frac{\pi}{360\,\arctan\left(\frac{0.5\,d_{\text{width}}}{r_{\text{w}}\,d_{\text{v}}}\right)},$ (1)
where $n_{\text{ppd}}$ is the effective display resolutions in pixels per degree, $d_{\text{width}}$ is the width of the screen, $r_{\text{w}}$ is the screen’s horizontal resolution in pixels and $d_{\text{v}}$ is the viewing distance. The display width and viewing distance must be provided in the same units (e.g. meters). We rely on $n_{\text{ppd}}$ later when expressing the spatial frequencies in cycles per visual degree. It should be noted that the above approximation is inaccurate for near-eye displays spanning a large field of view, and accurate geometric mapping should be used for these types of displays (see Equation 2 in Mantiuk et al. [2021]).
@@
[…]
$+L_{\text{refl}}+L_{\text{black}},L_{\text{peak}}\},$
@@
$L_{\text{refl}}=k_{\text{refl}}\frac{E_{\text{amb}}}{\pi}\,,$ (3)
@@
3.2 Opponent color channels
The sensitivity to chromatic changes is typically explained for color modulations represented in a space that separates three cardinal mechanisms of human color vision: achromatic channel and two chromatic channels, the latter commonly known as red-green and violet-yellow Stockman and Brainard [2010]. Here, we use the same color space as our contrast sensitivity function — the Derrington-Krauskopf-Lennie (DKL) colorspace Derrington et al. [1984]. The DKL space coordinates can be computed from the device-independent XYZ (provided by the display model) as:
$\begin{bmatrix}I_{\text{ach}}(\bm{x})\\ I_{\text{rg}}(\bm{x})\\ I_{\text{vy}}(\bm{x})\end{bmatrix}=\begin{bmatrix}1&1&0\\ 1&-\frac{L_{0}}{M_{0}}&0\\ -1&-1&\frac{L_{0}+M_{0}}{S_{0}}\end{bmatrix}M_{\text{XYZ}{\rightarrow}\text{LMS}}\begin{bmatrix}I_{\text{X}}(\bm{x})\\ I_{\text{Y}}(\bm{x})\\ I_{\text{Z}}(\bm{x})\end{bmatrix}\,,$ (4)
@@ 3.2 Opponent color channels
$M_{\text{XYZ}{\rightarrow}\text{LMS}}=\begin{bmatrix}0.187596&0.585169&-0.026384\\ -0.133397&0.405506&0.034502\\ 0.000244&-0.000543&0.019407\end{bmatrix}\,.$ (5)
@@
3.3 Temporal channels
Figure 3: The frequency characteristic of the four temporal channels used in ColorVideoVDP .
Psychophysical masking experiments showed evidence that the information is processed in the visual system by separate temporal channels. Two or three channels have been identified for the achromatic mechanism Anderson and Burr [1985]; Hess and Snowden [1992], and one or two channels for the chromatic mechanisms McKeefry et al. [2001]; Cass et al. [2009]. Here, we assume two temporal achromatic channels, as the third channel was observed only for low frequencies Hess and Snowden [1992]. We assume just one temporal channel for the red-green and violet-yellow cardinal directions as the evidence for the second channel shows that it has a less prominent role Cass et al. [2009]. Modeling fewer temporal channels also has the benefit of lower memory and computational cost.
The temporal channels of ColorVideoVDP  are directly defined by castleCSF Anonymous [2024] (see next section). The frequency tuning of the channels can be seen in Fig. 3. To find digital filters, we perform a real-valued (symmetric) fast Fourier transform on the frequency space filters. We found that a filter support of 250 ms is sufficient to capture filter characteristics. Finally, the two achromatic and two chromatic channels are convolved with the digital filters along the time dimension. This splits the achromatic signal into sustained (low-pass) and transient (band-pass) channels. The chromatic channels are low-pass filtered and become insensitive to high-frequency chromatic flicker.
3.4 Contrast sensitivity
Figure 4: castleCSF contrast sensitivity function for the four channels of ColorVideoVDP . The sensitivity is expressed in the cone-contrast units Wuerger et al. [2020]. Note that the achromatic transient and chromatic RG channels appear to have higher sensitivity than the achromatic sustained channel. This is due to the scaling used in the DKL color space Derrington et al. [1984] to represent chromatic contrast units. In practice, the transient and two chromatic channels are much less sensitive to patterns found in complex images.
A contrast sensitivity function models our ability to detect patterns of different spatial and temporal frequency, size, and shown at different luminance levels. It is a cornerstone of ColorVideoVDP  — it defines the temporal and chromatic channels and enables modeling of contrast masking, which is the key component of our metric. Since there is no contrast sensitivity model that could explain sensitivity to both chromatic modulations and different temporal frequencies, we have created our own model, named castleCSF. This CSF models color, area, spatial and temporal frequencies, luminance, and eccentricity. Because of its complexity, castleCSF is explained in detail in a separate paper Anonymous [2024]. Here, we summarize its key components.
@@ 3.4 Contrast sensitivity
To find the detection threshold, castleCSF modulates (multiplies) the contrast associated with each mechanism by the sensitivity of that mechanism and then pools those to form the contrast energy. The detection threshold is assumed to be the contrast at which the energy is equal to 1. The sensitivity of each mechanism is modeled as a function of spatial and temporal frequency, area, background luminance and eccentricity, in a similar manner as for stelaCSF Mantiuk et al. [2022]. castleCSF was optimized to predict the data from 19 contrast sensitivity datasets (10 achromatic, 6 chromatic and 3 mixed).
@@
3.5 Multi-scale decomposition and color contrast
In addition to the temporally-tuned channels, psychophysical Foley [1994]; Stromeyer and Julesz [1972] and neuropsychological De Valois et al. [1982] evidence proves the existence of channels that are tuned to the bands of spatial frequencies and orientations. We must account for such channels to model visual masking, as explained in the next section. Following FovVideoVDP Mantiuk et al. [2021], we use the Laplacian pyramid Burt and Adelson [1983] to decompose each of the four temporal channels (Y-sustained, Y-transient, RG, YV) into spatial-frequency selective bands. However, unlike FovVideoVDP, we consider low frequencies, including the base band (the coarsest low-pass band in the Laplacian pyramid). This change is due to finding that many display distortions can only be detected in the low-spatial frequency bands (see Fig. 1). We select the height of the pyramid so that the lowest peak frequency of the band-pass filter is greater than 0.2 cpd. Similarly as in FovVideoVDP, we do not consider orientation-selective channels because of the substantial computational overhead of those.
@@ 3.5 Multi-scale decomposition and color contrast
$C_{b,c,f}(\bm{x})=\frac{\mathcal{L}_{b,c,f}(\bm{x})}{\Uparrow\mathcal{G}_{b+1,\mathcal{S},f}(\bm{x})}=\frac{\mathcal{L}_{b,c,f}(\bm{x})}{L_{\text{bkg},b,f}(\bm{x})}\,,$ (6)
where $\mathcal{L}_{b,c,f}$ represents the $b$-th band of the Laplacian pyramid, $\mathcal{G}_{b+1,\mathcal{S},f}$ is the $(b+1)$ band of the Gaussian pyramid and $\Uparrow$ is the upsampling (expand) operator. The Laplacian pyramid is created for each frame of each channel, obtained by temporal filtering (Sec. 3.3). The Laplacian pyramid coefficients are divided by the values from the Gaussian pyramid at the band $b+1$ (one lever coarser) of the sustained luminance channel ($\mathcal{S}$). Note that the upsampled version of the Gaussian pyramid band $b+1$ is a by-product of computing a Laplacian pyramid so can be obtained at no computational cost. This formulation is similar to that of local band-limited contrast Peli [1990], but here we extend it to both achromatic and chromatic channels. These contrast values are computed separately for the test and reference frames. The values in all bands except the low-pass (base-band) and the high-pass bands are multiplied by 2 to account for the reduction in amplitude [Mantiuk et al., 2021, Fig. 5].
3.6 Cross-channel contrast masking
@@
Figure 5: Example of contrast masking. A sinusoidal grating of 4 cpd (when seen printed from 40 cm) was added to a reference image (top row) to obtain a distorted image (second row). The image was split into three parts and the grating was modulated along achromatic, red-green, and yellow-violet directions in each respective part. The third row shows the visual difference map generated by ColorVideoVDP  without masking but still using the CSF. The map over-predicts the visibility in textured areas. The bottom row shows the prediction with the masking model. It is worth noting: although all three color directions are equally visible in the gray bar at the bottom of the image (where there is little masking), the red-green pattern is more visible in the textured area because it is weakly masked by the achromatic channel. ColorVideoVDP  correctly predicts this phenomenon.
@@
The model of cross-channel contrast masking transforms local physical contrast differences between the test and reference frames into perceived differences — differences that are scaled by the local contrast visibility. It accounts for both contrast sensitivity, such as lower sensitivity to high spatial and temporal frequencies, and contrast masking. Masking accounts for differences being less likely to be noticed in heavily textured areas. The cross-channel component models that a strong contrast in one channel can reduce the visibility of contrast in another channel Switkes et al. [1988], for example, contrast in the chromatic channels can mask contrast in achromatic channels. Finally, the masking model also needs to account for suprathreshold contrast perception — must match the perceived magnitude of contrast (changes) across luminance, frequencies, and the directions of color modulation.
Contrast masking is the critical component of the metric that determines its performance, as we show later in the ablation studies (Sec. 5.2). To make the optimal choice, we analyzed and compared six models: two contrast encodings (additive and multiplicative) combined with three masking models. We tested the mutual masking model from the original VDP Daly [1993], the contrast transducer proposed by Watson and Solomon [1997], and the contrast similarity formula used in SSIM and many other metrics. To keep this text concise, we describe below only the mutual masking model, which performed the best in our tests. We encourage the reader to refer to the supplementary materials with the detailed description and analysis of all the models.
Contrast masking shows different characteristics for different spatial frequency bands and color channels. However, it was shown that for both luminance Daly [1993] and chromatic channels Cass et al. [2009]; Switkes et al. [1988] masking characteristics can be unified if both the test and masker contrast are multiplied by the contrast sensitivity function $S_{b,c}(\bm{x})$:
$C^{\prime}_{b,c,f}(\bm{x})=C_{b,c,f}(\bm{x})\,S_{b,c,f}(\bm{x})\,.$ (7)
@@
$S_{b,c,f}(\bm{x})=s_{\text{corr}}\,s_{\text{ch},c}\,S_{c}(\rho_{b},\omega_{c},L_{\text{bkg},b,f}^{\text{ref}}(\bm{x}))\,,$ (8)
@@
$D_{b,c,f}(\bm{x})=\frac{\left|C^{\prime\,\textrm{test}}_{b,c,f}(\bm{x})-C^{\prime\,\textrm{ref}}_{b,c,f}(\bm{x})\right|^{p}}{1+C^{\textrm{mask}}_{b,c,f}(\bm{x})}\,,$ (9)
where $p$ is a parameter of the model. The masking signal $C^{\textrm{mask}}_{b,c}$ combines local contrast across test and reference images, local spatial neighborhood, and channels. First, similarly as in [Daly, 1993, p.192], we calculate the mutual masking of test and reference bands (see also Fig. 6):
$C^{\text{mm}}_{b,c,f}(\bm{x})=\min\left\{\left|C^{\prime\,\textrm{test}}_{b,c,f}(\bm{x})\right|,\left|C^{\prime,\textrm{ref}}_{b,c,f}(\bm{x})\right|\right\}\,.$ (10)
@@
$C^{\textrm{mask}}_{b,c,f}(\bm{x})=\sum_{i}k_{i,c}((C^{\text{mm}}_{b,i,f})^{q_{c}}*g_{\sigma_{\text{sp}}})(\bm{x})\,.$ (11)
where $k_{i,c}$ is the cross-channel masking coefficient, describing the contribution of channel $i$ to the masking signal of channel $c$. The weights for the trained model are shown in Fig. 9. The exponents $q_{c}$ are model parameters set separately for each channel (two achromatic and two chromatic channels).
@@
$D_{B,c,f}(\bm{x})=k_{\text{B},c}\,\left|\mathcal{G}_{B,c,f}^{\text{test}}(\bm{x})-\mathcal{G}_{B,c,f}^{\text{ref}}(\bm{x})\right|\,S_{B,c,f}(\bm{x})\,,$ (12)
@@
$\hat{D}_{B,c,f}(\bm{x})=\frac{k_{\text{C}}\,D_{B,c,f}(\bm{x})}{k_{\text{C}}+D_{B,c,f}(\bm{x})}\,,$ (13)
@@
The shape of the masking function for the sustained and transient achromatic channels is plotted in Fig. 7. Another visualization of the masking model is shown Fig. 8 as a discrimination vs. masking plot. The plot shows that, as the mutual masking contrast $C^{\text{mm}}$ increases, a higher difference between the test and reference bands is needed to trigger the same response ($D_{b,c,f}(\bm{x})=1$ in this example). Such an increase is more gradual for the sustained achromatic channel and chromatic channels and more abrupt for the transient channel. The findings of Switkes et al. Switkes et al. [1988] indicate that color can robustly mask luminance patterns, while luminance does not mask color but instead facilitates the detection of color patterns. Remarkably, the cross-channel masking weights obtained in metric fitting (Sec. 5) and shown in Fig. 9 confirm these findings — weights are very small in the 2nd and 3rd column of the top row, showing little influence of luminance (achromatic sustained channel) on masking of the color channels. Note that our mutual masking model cannot account for facilitation. An example of this effect is shown in Fig. 5, in which luminance contrast does not mask red-green chromatic contrast.
Figure 9: Weights of the cross-channel masking model — $k_{i,c}$. The values are normalized to one in each column to better show the contribution of each channel. The absolute values can be obtained by multiplying the cell value by a constant shown in the parenthesis at the bottom of each column. The matrix shows that both chromatic channels strongly mask the sustained achromatic channel (1st column), but the achromatic channels do not mask chromatic channels.
3.7 Pooling
Once the differences between the bands, $D_{b,c,f}$, are computed, they need to be pooled into a single quality correlate. We follow a similar strategy as done in FovVideoVDP and pool the differences across all the spatial dimensions ($\bm{x}$) in each band, across spatial frequency bands ($b$), across the channels ($c$) and finally across all the frames ($f$):
$D_{\textrm{pooled}}=\frac{1}{F^{\nicefrac{{1}}{{\beta_{f}}}}}\left\|\left\|w_{c}\left\|\frac{1}{N_{b}^{\nicefrac{{1}}{{\beta_{x}}}}}\left\|\hat{D}_{b,c,f}(\bm{x})\right\|_{\beta_{x},\bm{x}}\right\|_{\beta_{b},b}\right\|_{\beta_{c},c}\right\|_{\beta_{f},f}\,,$ (14)

### v2

@@ 3.1. Display model
(1) $n_{\text{ppd}}=\frac{\pi}{360\,\arctan\left(\frac{0.5\,d_{\text{width}}}{r_{\text{w}}\,d_{\text{v}}}\right)},$
where $n_{\text{ppd}}$ is the effective display resolutions in pixels per degree, $d_{\text{width}}$ is the width of the screen, $r_{\text{w}}$ is the screen’s horizontal resolution in pixels and $d_{\text{v}}$ is the viewing distance. The display width and viewing distance must be provided in the same units (e.g. meters). We rely on $n_{\text{ppd}}$ later when expressing the spatial frequencies in cycles per visual degree. It should be noted that the above approximation is inaccurate for near-eye displays spanning a large field of view, and accurate geometric mapping should be used for these types of displays (see Equation 2 in (Mantiuk et al., 2021)).
@@
(2) $I_{\text{lin},c}(\bm{x})=\min\left\{(L_{\text{peak}}-L_{\text{black}})\,E(I_{\text{de},c}(\bm{x}))+L_{\text{black}},L_{\text{peak}}\right\}+L_{\text{refl}},$
@@
(3) $L_{\text{refl}}=k_{\text{refl}}\frac{E_{\text{amb}}}{\pi}\,,$
@@
3.2. Opponent color channels
The sensitivity to chromatic changes is typically explained for color modulations represented in a space that separates three cardinal mechanisms of human color vision: achromatic channel and two chromatic channels, the latter commonly known as red-green and violet-yellow (Stockman and Brainard, 2010). Here, we use the same color space as our contrast sensitivity function — the Derrington-Krauskopf-Lennie (DKL) colorspace (Derrington et al., 1984). DKL was selected for castleCSF as this color space is linearly related to cone responses, is well established in vision science, and much of the chromatic contrast detection data was collected in that space. The DKL space coordinates can be computed from the device-independent XYZ (provided by the display model) as:
(4) $\begin{bmatrix}I_{\text{ach}}(\bm{x})\\ I_{\text{rg}}(\bm{x})\\ I_{\text{vy}}(\bm{x})\end{bmatrix}=\begin{bmatrix}1&1&0\\ 1&-\frac{L_{0}}{M_{0}}&0\\ -1&-1&\frac{L_{0}+M_{0}}{S_{0}}\end{bmatrix}M_{\text{XYZ}{\rightarrow}\text{LMS}}\begin{bmatrix}I_{\text{X}}(\bm{x})\\ I_{\text{Y}}(\bm{x})\\ I_{\text{Z}}(\bm{x})\end{bmatrix}\,,$
@@ 3.2. Opponent color channels
(5) $M_{\text{XYZ}{\rightarrow}\text{LMS}}=\begin{bmatrix}0.187596&0.585169&-0.026384\\ -0.133397&0.405506&0.034502\\ 0.000244&-0.000543&0.019407\end{bmatrix}\,.$
@@
3.3. Temporal channels
Figure 3. The frequency characteristic of the four temporal channels used in ColorVideoVDP.
Psychophysical masking experiments showed evidence that the information is processed in the visual system by separate temporal channels. Two or three channels have been identified for the achromatic mechanism (Anderson and Burr, 1985; Hess and Snowden, 1992), and one or two channels for the chromatic mechanisms (McKeefry et al., 2001; Cass et al., 2009). Here, we assume two temporal achromatic channels, as the third channel was observed only for low frequencies (Hess and Snowden, 1992). We assume just one temporal channel for the red-green and violet-yellow cardinal directions as the evidence for the second channel shows that it has a less prominent role (Cass et al., 2009). Modeling fewer temporal channels also has the benefit of lower memory and computational cost.
The temporal channels of ColorVideoVDP are directly defined by castleCSF (Ashraf et al., 2024) (see next section). The frequency tuning of the channels can be seen in Fig. 3. To find digital filters, we perform a real-valued (symmetric) fast Fourier transform on the frequency space filters. We found that a filter support of 250 ms is sufficient to capture filter characteristics. Finally, the two achromatic and two chromatic channels are convolved with the digital filters along the time dimension. This splits the achromatic signal into sustained (low-pass) and transient (band-pass) channels. The chromatic channels are low-pass filtered and become insensitive to high-frequency chromatic flicker.
3.4. Contrast sensitivity
Figure 4. castleCSF contrast sensitivity function for the four channels of ColorVideoVDP. The sensitivity is expressed in the cone-contrast units (Wuerger et al., 2020). Note that the achromatic transient and chromatic RG channels appear to have higher sensitivity than the achromatic sustained channel. This is due to the scaling used in the DKL color space (Derrington et al., 1984) to represent chromatic contrast units. In practice, the transient and two chromatic channels are much less sensitive to patterns found in complex images.
A contrast sensitivity function models our ability to detect patterns of different spatial and temporal frequency, size, and shown at different luminance levels. It is a cornerstone of ColorVideoVDP — it defines the temporal and chromatic channels and enables modeling of contrast masking, which is the key component of our metric. Since there is no contrast sensitivity model that could explain sensitivity to both chromatic modulations and different temporal frequencies, we have created our own model, named castleCSF. This CSF models color, area, spatial and temporal frequencies, luminance, and eccentricity. Because of its complexity, castleCSF is explained in detail in a separate paper (Ashraf et al., 2024). Here, we summarize its key components.
@@ 3.4. Contrast sensitivity
To find the detection threshold, castleCSF modulates (multiplies) the contrast associated with each mechanism by the sensitivity of that mechanism and then pools those to form the contrast energy. The detection threshold is assumed to be the contrast at which the energy is equal to 1. The sensitivity of each mechanism is modeled as a function of spatial and temporal frequency, area, background luminance and eccentricity, in a similar manner as for stelaCSF (Mantiuk et al., 2022). castleCSF was optimized to predict the data from 19 contrast sensitivity datasets (10 achromatic, 6 chromatic and 3 mixed).
@@
3.5. Multi-scale decomposition and color contrast
In addition to the temporally-tuned channels, psychophysical (Foley, 1994; Stromeyer and Julesz, 1972) and neuropsychological (De Valois et al., 1982) evidence proves the existence of channels that are tuned to the bands of spatial frequencies and orientations. We must account for such channels to model visual masking, as explained in the next section. Following FovVideoVDP (Mantiuk et al., 2021), we use the Laplacian pyramid (Burt and Adelson, 1983) to decompose each of the four temporal channels (Y-sustained, Y-transient, RG, YV) into spatial-frequency selective bands. However, unlike FovVideoVDP, we consider low frequencies, including the base band (the coarsest low-pass band in the Laplacian pyramid). This change is due to finding that many display distortions can only be detected in the low-spatial frequency bands (see Fig. 1). We select the height of the pyramid so that the lowest peak frequency of the band-pass filter is greater than 0.2 cpd. Similarly as in FovVideoVDP, we do not consider orientation-selective channels because of the substantial computational overhead of those.
@@ 3.5. Multi-scale decomposition and color contrast
(6) $C_{b,c,f}(\bm{x})=\frac{\mathcal{L}_{b,c,f}(\bm{x})}{\Uparrow\mathcal{G}_{b+1,\mathcal{S},f}(\bm{x})}=\frac{\mathcal{L}_{b,c,f}(\bm{x})}{L_{\text{bkg},b,f}(\bm{x})}\,,$
where $\mathcal{L}_{b,c,f}$ represents the $b$-th band of the Laplacian pyramid, $\mathcal{G}_{b+1,\mathcal{S},f}$ is the $(b+1)$ band of the Gaussian pyramid and $\Uparrow$ is the upsampling (expand) operator. The Laplacian pyramid is created for each frame of each channel, obtained by temporal filtering (Sec. 3.3). The Laplacian pyramid coefficients are divided by the values from the Gaussian pyramid at the band $b+1$ (one lever coarser) of the sustained luminance channel ($\mathcal{S}$). Note that the upsampled version of the Gaussian pyramid band $b+1$ is a by-product of computing a Laplacian pyramid so can be obtained at no computational cost. This formulation is similar to that of local band-limited contrast (Peli, 1990), but here we extend it to both achromatic and chromatic channels. These contrast values are computed separately for the test and reference frames. The values in all bands except the low-pass (base-band) and the high-pass bands are multiplied by 2 to account for the reduction in amplitude (Mantiuk et al., 2021, Fig. 5).
3.6. Cross-channel contrast masking
@@
Figure 5. Example of contrast masking. A sinusoidal grating of 4 cpd (when seen printed from 40 cm) was added to a reference image (top row) to obtain a distorted image (second row). The image was split into three parts and the grating was modulated along achromatic, red-green, and yellow-violet directions in each respective part. The third row shows the visual difference map generated by ColorVideoVDP without masking but still using the CSF. The map over-predicts the visibility in textured areas. The bottom row shows the prediction with the masking model. It is worth noting: although all three color directions are equally visible in the gray bar at the bottom of the image (where there is little masking), the red-green pattern is more visible in the textured area because it is weakly masked by the achromatic channel. ColorVideoVDP correctly predicts this phenomenon.
@@
The model of cross-channel contrast masking transforms local physical contrast differences between the test and reference frames into perceived differences — differences that are scaled by the local contrast visibility. It accounts for both contrast sensitivity, such as lower sensitivity to high spatial and temporal frequencies, and contrast masking. Masking accounts for differences being less likely to be noticed in heavily textured areas. The cross-channel component models that a strong contrast in one channel can reduce the visibility of contrast in another channel (Switkes et al., 1988), for example, contrast in the chromatic channels can mask contrast in achromatic channels. Finally, the masking model also needs to account for suprathreshold contrast perception — must match the perceived magnitude of contrast (changes) across luminance, frequencies, and the directions of color modulation.
Contrast masking is the critical component of the metric that determines its performance, as we show later in the ablation studies (Sec. 5.2). To make the optimal choice, we analyzed and compared six models: two contrast encodings (additive and multiplicative) combined with three masking models. We tested the mutual masking model from the original VDP (Daly, 1993), the contrast transducer proposed by Watson and Solomon (1997), and the contrast similarity formula used in SSIM and many other metrics. To keep this text concise, we describe below only the mutual masking model, which performed the best in our tests. We encourage the reader to refer to the supplementary materials with the detailed description and analysis of all the models.
Contrast masking shows different characteristics for different spatial frequency bands and color channels. However, it was shown that for both luminance (Daly, 1993) and chromatic channels (Cass et al., 2009; Switkes et al., 1988) masking characteristics can be unified if both the test and masker contrast are multiplied by the contrast sensitivity function $S_{b,c}(\bm{x})$:
(7) $C^{\prime}_{b,c,f}(\bm{x})=C_{b,c,f}(\bm{x})\,S_{b,c,f}(\bm{x})\,.$
@@
(8) $S_{b,c,f}(\bm{x})=s_{\text{corr}}\,s_{\text{ch},c}\,S_{c}(\rho_{b},\omega_{c},L_{\text{bkg},b,f}^{\text{ref}}(\bm{x}))\,,$
@@
(9) $D_{b,c,f}(\bm{x})=\frac{\left|C^{\prime\,\textrm{test}}_{b,c,f}(\bm{x})-C^{\prime\,\textrm{ref}}_{b,c,f}(\bm{x})\right|^{p}}{1+C^{\textrm{mask}}_{b,c,f}(\bm{x})}\,,$
where $p$ is a parameter of the model. The masking signal $C^{\textrm{mask}}_{b,c}$ combines local contrast across test and reference images, local spatial neighborhood, and channels. First, similarly as in (Daly, 1993, p.192), we calculate the mutual masking of test and reference bands (see also Fig. 6):
(10) $C^{\text{mm}}_{b,c,f}(\bm{x})=\min\left\{\left|C^{\prime\,\textrm{test}}_{b,c,f}(\bm{x})\right|,\left|C^{\prime,\textrm{ref}}_{b,c,f}(\bm{x})\right|\right\}\,.$
@@
(11) $C^{\textrm{mask}}_{b,c,f}(\bm{x})=\sum_{i}k_{i,c}((C^{\text{mm}}_{b,i,f})^{q_{c}}*g_{\sigma_{\text{sp}}})(\bm{x})\,.$
@@
[…]
(12) $D_{B,c,f}(\bm{x})=k_{\text{B},c}\,\left|\mathcal{G}_{B,c,f}^{\text{test}}(\bm{x})-\mathcal{G}_{B,c,f}^{\text{ref}}(\bm{x})\right|\,S_{B,c,f}(\bm{x})\,,$
@@
(13) $\hat{D}_{B,c,f}(\bm{x})=\frac{k_{\text{C}}\,D_{B,c,f}(\bm{x})}{k_{\text{C}}+D_{B,c,f}(\bm{x})}\,,$
@@
The shape of the masking function for the sustained and transient achromatic channels is plotted in Fig. 7. Another visualization of the masking model is shown Fig. 8 as a discrimination vs. masking plot. The plot shows that, as the mutual masking contrast $C^{\text{mm}}$ increases, a higher difference between the test and reference bands is needed to trigger the same response ($D_{b,c,f}(\bm{x})=1$ in this example). Such an increase is more gradual for the sustained achromatic channel and chromatic channels and more abrupt for the transient channel. The findings of Switkes et al. (1988) indicate that color can robustly mask luminance patterns, while luminance does not mask color but instead facilitates the detection of color patterns. Remarkably, the cross-channel masking weights obtained in metric fitting (Sec. 5) and shown in Fig. 9 confirm these findings — weights are very small in the 2nd and 3rd column of the top row, showing little influence of luminance (achromatic sustained channel) on masking of the color channels. Note that our mutual masking model cannot account for facilitation. An example of this effect is shown in Fig. 5, in which luminance contrast does not mask red-green chromatic contrast.
Figure 9. Weights of the cross-channel masking model — $k_{i,c}$. The values are normalized to one in each column to better show the contribution of each channel. The absolute values can be obtained by multiplying the cell value by a constant shown in the parenthesis at the bottom of each column. The matrix shows that both chromatic channels strongly mask the sustained achromatic channel (1st column), but the achromatic channels do not mask chromatic channels.
3.7. Pooling
Once the differences between the bands, $D_{b,c,f}$, are computed, they need to be pooled into a single quality correlate. We follow a similar strategy as done in FovVideoVDP and pool the differences across all the spatial dimensions ($\bm{x}$) in each band, across spatial frequency bands ($b$), across the channels ($c$), and finally across all the frames ($f$):
(14) $D_{\textrm{pooled}}=\frac{1}{F^{\nicefrac{{1}}{{\beta_{f}}}}}\left\|\left\|w_{c}\left\|\frac{1}{N_{b}^{\nicefrac{{1}}{{\beta_{x}}}}}\left\|\hat{D}_{b,c,f}(\bm{x})\right\|_{\beta_{x},\bm{x}}\right\|_{\beta_{b},b}\right\|_{\beta_{c},c}\right\|_{\beta_{f},f}\,,$